OpenStudy (anonymous):

use power rules to simplify (2xy^6)^5

OpenStudy (zepp):

Power rule:\[\LARGE y=x^n\\\LARGE y'=nx^{n-1}\]

OpenStudy (anonymous):

\[\large (2xy^6)^5 = (2)^5 \cdot (x)^5 \cdot (y^6)^5\]

Parth (parthkohli):

Why do we have to make it so complicated?

OpenStudy (zepp):

Power of a power rule:\[\LARGE (x^a)^b=x^{a*b}\]

OpenStudy (lgbasallote):

i think by power rule she meant \[\huge (ab)^m = a^m b^m\]

Parth (parthkohli):

Distribute 5 into all the exponents.

OpenStudy (anonymous):

\[\huge (y^a)^b = y^{ a \times b}\]

Parth (parthkohli):

\((xy)^z = x^z \times y^z\)

OpenStudy (lgbasallote):

@legs3639 http://openstudy.com/updates/5014bfcce4b0fa24673114a7

OpenStudy (zepp):

People, stop flooding/spamming and the most important.. stop repeating what I've typed .. .-.

OpenStudy (anonymous):

im still at a loss here

OpenStudy (zepp):

And I'm pretty sure the asker could easily figure out what to do with the only two hints I typed.

OpenStudy (anonymous):

i wish that was true

OpenStudy (lgbasallote):

check the link i posted...i gave an explanation to it already

OpenStudy (anonymous):

ty

OpenStudy (zepp):

The first thing to do is to simplify it, then you take the derivative with the power rule.

OpenStudy (anonymous):

ty

OpenStudy (lgbasallote):

@zepp this is an algebra question...not a calculus one

OpenStudy (zepp):

Oh

OpenStudy (zepp):

Power rule sounds pretty Calculus to me.

OpenStudy (lgbasallote):

and for yur question why the explanations are being repeated..it's because everyone's so competitive